[figure: 9.7 cm]

TEMPERATURE DEPENDENCE OF THE INTERNAL ENERGY

FOR ISING MODEL ±J ON THE TUBE

Marian Błaszyk Institute of Physics, A. Mickiewicz University, Poznań, Poland

A method of linear equations has been applied for determination of the temperature dependence of internal energy. The method permits a reduction of a physical problem of determination of the temperature dependence of internal energy for systems with frozen disorder to a mathematical question of solving a set of linear equations. The Ising model with exchange integrals ±J on a square lattice spanned over a tube of infinite length and the width of 2 bonds is considered.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : M. Błaszyk

Address for correspondence : Institute of Physics, A. Mickiewicz University, Poznań, ul. Umultowska 85, Poland

Email address : blaszyk@main.amu.edu.pl